SHORT COMMUNICATIONS

D. Apushkinskaya: The initial-boundary value problem to nondivergent parabolic equations with the Vent’cel boundary condition F. Brock: Steiner symmetrization and periodic solutions of boundary and initial value problems V.I. Burenkov, M.L. Gol’dman, L.-E. Persson, V.D. Stepanov: Best constants in weight inequalities for positive operators on the cone of monotone functions J. García–Cuerva: Vector-valued Hausdorff–Young inequalities C.E. Gutierrez: Convex sets related to the Monge–Ampère equations A. Ilyin: Averaging principle for dissipative PSEs E.A. Kalita: Dual Morrey spaces and singular solvability of nonlinear elliptic systems V. Kokilashvili: Improvement of classical imbedding theorems in scale of weighted function spaces E.R. Love: Expansions in series of Legendre functions R. Oinarov: Three-weights inequality for norms of intermediate derivatives S. Pérez: Endpoint estimates for commutators of singular integrals A. Stanoyevitch: Steiner symmetrization and Poincaré inequalities J.Q. Sun: Two weight mixed-Φ-inequalities for operators of Hardy type